# GraphInsight: Unlocking Insights in Large Language Models for Graph Structure Understanding

### v2

@@ 5.1 Experimental Setup
We compare GraphInsight with 9 baselines: (a) raw sequence input method w/o any processing (Raw Seq.), (b) prompting methods including Build-a-Graph Prompting (BAG) Wang et al. (2024a), Chain-of-Thought (COT) Wei et al. (2022) and Few-Shot (FS) Brown et al. (2020), (c) reordering methods Ge et al. (2024) based on breadth-first-search order (BFS), depth-first-search order (DFS), and shortest-path order (SP), (d) structural methods including adjacency lists (AL), and adjacency matrices (AM).
@@
Existing graph understanding benchmarks suffer from limited node coverage, unclear task definitions, a lack of structural diversity, and insufficient support for multi-step graph reasoning. To address these issues, we introduce GraphSQA, a comprehensive benchmark designed to evaluate the capabilities of LLMs in understanding graph structures, which encompasses a wide range of task types found in existing benchmarks Wang et al. (2024a); Yuan et al. (2024). GraphSQA includes a broad spectrum of node counts and features diverse graph structures, such as multi-edges and self-loops. It encompasses two types of graph understanding tasks: (1) 5 macro-level graph tasks and (2) 15 micro graph tasks including 8 composite tasks. Further details are provided in Appendix Section 1. In our experiments, we evaluate the average performance of various methods and models on diverse graph tasks with $|V|$ ranging from 15 to 200 nodes, covering 10,400 tasks in total—2,600 macro-level and 7,800 micro-level.
@@
In GraphSQA, there are three types of answers: boolean, numerical, and set types. For boolean answers, the score is 1 if the answer matches the ground truth; otherwise, the score is 0. For numerical answers, we use one minus relative error as the score Wang et al. (2024a). For answers of set type, we use the Jaccard similarity Ji et al. (2013) between the answer set and the ground truth set as the score.
@@
We utilized a diverse set of open source models, including those fine-tuned for long sequences including Mistral 7B (small-scale) and Llama3-8B-256K (fine-tuned for long sequences), and their non-long-sequence counterparts including Llama3 8B, Qwen2-7B-Instruct, and Vicuna-7B.
COT FS BAG BFS DFS SP AL AM
Mistral-7B Overall 0.5476 0.5315 0.4671 0.5145 0.5481 0.5482 0.5486 0.5320 0.1922 0.6811
Macro 0.5222 0.4323 0.5439 0.4249 0.5212 0.5205 0.5222 0.5001 0.1808 0.5846
Micro 0.5635 0.5935 0.4192 0.5705 0.5650 0.5655 0.5651 0.5519 0.1994 0.7425
Llama-3-8B Overall 0.4513 0.4504 0.4173 0.4358 0.4551 0.4591 0.4559 0.4048 0.1434 0.6765
Macro 0.4379 0.4187 0.3702 0.4060 0.4381 0.4376 0.4404 0.4336 0.1198 0.5422
Micro 0.4597 0.4702 0.4468 0.4545 0.4658 0.4726 0.4656 0.3869 0.1581 0.7605
Qwen2-7B Overall 0.5640 0.5663 0.5400 0.5677 0.5381 0.5290 0.5292 0.5595 0.2129 0.7695
Macro 0.5644 0.5423 0.5366 0.6447 0.6017 0.5913 0.5958 0.5223 0.1915 0.6587
Micro 0.5637 0.5813 0.5421 0.5196 0.4984 0.4901 0.4876 0.5828 0.2262 0.8387
Llama-3-8B-262k Overall 0.7183 0.5279 0.5825 0.6231 0.7150 0.7166 0.7147 0.6603 0.5038 0.8285
Macro 0.6218 0.5751 0.4652 0.5115 0.6189 0.6221 0.6183 0.5292 0.3529 0.6928
Micro 0.7786 0.4985 0.6559 0.6928 0.7752 0.7757 0.7750 0.7422 0.5982 0.9401
Vicuna-7B Overall 0.1273 0.1147 0.1041 0.1066 0.1276 0.1277 0.1277 0.0862 0.0345 0.3557
Macro 0.1267 0.1164 0.0877 0.0995 0.1266 0.1259 0.1263 0.0567 0.0409 0.2296
Micro 0.1276 0.1137 0.1144 0.1111 0.1282 0.1289 0.1285 0.1046 0.0306 0.4346
@@ 5.2 Performance
GraphInsight outperforms all other methods across all large models for macro-tasks, as shown in Table 1. Specifically, GraphInsight can achieve up to a 4.61$\times$ increase in score compared to AM on the Vicuna-7B model. Structural methods show the smallest improvement for macro-level tasks. The other two types of methods, i.e., prompting and structural methods, result in only minimal improvements in the understanding of LLMs for macro-level tasks. For example, on Llama-3-8B, they can achieve at most a minimal improvement of 5.7%, from 0.4379 to 0.4404. However, GraphInsight can provide a substantial increase of 23.82%, from 0.4379 to 0.5422. Also, as shown in Figures 4(a)-(b), across macro-level tasks with varying $|V|$, GraphInsight consistently outperforms other methods. As $|V|$ increases, the understanding capability tends to decrease, but GraphInsight shows the smallest decline.
@@
(a) Macro-level (Mistral-7B) (b) Macro-level (Vicuna-7B) Refer to caption Refer to caption (c) Micro-level (Mistral-7B) (d) Micro-level (Vicuna-7B)
Figure 4: Analysis on Graphs with Different $|V|$
(a) Qwen2-7B (b) Vicuna-7B
Figure 5: Analysis on Composite Tasks
@@ 5.3 Ablation Study
We conduct the ablation study on graph description organization and GraphRAG to verify the effectiveness of them. As shown in Table 2, graph description organization is the optimization primarily for macro-level tasks as well as for micro-level tasks. For example, on Llama-3-8B-262K, reorganization can achieve an 11.4% increase for macro-level tasks and an 8.9% increase for micro-level tasks. On the other hand, GraphRAG focuses on micro-level tasks and can provide further notably improvements for micro-level tasks on top of the gains achieved by reorganization. For example, GraphRAG brings a 173.7% improvement on Vicuna-7B.
@@ 5.4 Hyperparameter Analysis
We conduct two types of experiments involving $\alpha\%$ and $\beta\%$. We fix $\alpha\%+\beta\%=15\%$ (resp. fix $\frac{\alpha\%}{\beta\%}=\frac{3}{7}$) to examine how variations in $\frac{\alpha\%}{\beta\%}$ (resp. $\alpha\%+\beta\%$) impact the graph understanding of LLMs. As shown in Figure 6, the LLMs’ understanding remains relatively stable in both scenarios, indicating insensitivity to these parameter changes. For simplicity, we recommend setting $\frac{\alpha\%}{\beta\%}$ to $\frac{3}{7}$ and $\alpha\%+\beta\%$ to 15%.
@@
The variation of $\gamma$affects the size of the GraphRAG base. As $\gamma$increase, the GraphRAG base can retain more important graph structural information, leading to better micro-level understanding of LLMs. It can be observed from Figure 6 that when $\gamma$exceeds 80%, the improvement in the model’s understanding becomes less significant. We recommend setting $\gamma$to 80%.
(c) Mistral-7B ($\alpha+\beta=15$) (d) Qwen2-7B ($\alpha+\beta=15$)
Figure 6: Head and Tail Strong Region Thresholds $\alpha$ and $\beta$.
(a) Mistral-7B (b) Qwen2-7B
Figure 7: GraphRAG Rate $\gamma\%$.
@@ 6 Conclusion
In this paper, we introduce GraphInsight, the first framework that analyzes and addresses the challenge of graph structure comprehension by LLMs from the perspective of their inherent “positional bias.” By leveraging LLMs’ strong memory in certain graph sequence regions and compensating for weaker ones, GraphInsight enhances LLMs’ graph understanding. The framework employs two key techniques: importance-based graph description reorganization and lightweight RAG, tailored for macro- and micro-level graph understanding tasks. Extensive experiments show that GraphInsight outperforms all baselines in comprehending graph structures of various sizes. In the future, we aim to improve LLMs’ understanding of more complex graph types, including those with labels and semantic information.
Appendix
@@ A.1 Introduction to GraphSQA
GraphSQA offers a standardized evaluation suite, enabling consistent and comprehensive assessment of LLMs’ capabilities in understanding and reasoning about graph structures. Unlike existing benchmarks that primarily focus on small graphs containing within 35 nodes, GraphSQA broadens the scope by encompassing a diverse spectrum of node quantities, ranging from 15 to 300 nodes, and features an array of graph structures, including multi-edges and self-loops. The benchmark is organized into two primary categories of graph understanding tasks: macro-level tasks and micro-level tasks, with the latter further subdivided into 15 tasks, including 8 composite tasks, which provides a more challenging and extensive evaluation environment, better reflecting the complexities encountered in real-world graph-based applications.
@@ Appendix D Notations
Appendix E Metric Analysis
@@
Appendix F Performance Analysis of the GraphInsight Framework
@@ Appendix G Discussion
Appendix H Task Templates in GraphSQA

### v3

@@ 5.1 Experimental Setup
We compare GraphInsight with 10 baselines: (a) raw sequence input method without any processing (Raw Seq.), (b) GNN encoding-based method [^3] GNN encoding-based methods like GraphToken rely on implicit graph embeddings, which are orthogonal to GraphInsight’s explicit graph description approach. Despite this, we include them to ensure a comprehensive comparison. (GraphToken) Perozzi et al. (2024), (c) prompting methods including Build-a-Graph Prompting (BAG) Wang et al. (2024a), Chain-of-Thought (COT) Wei et al. (2022), and Few-Shot (FS) Brown et al. (2020), (d) reordering methods Ge et al. (2024) based on breadth-first-search order (BFS), depth-first-search order (DFS), and shortest-path order (SP), (e) structural methods including adjacency lists (AL) and adjacency matrices (AM).
(a) Macro-level (Mistral-7B) (b) Macro-level (Vicuna-7B)
(c) Micro-level (Mistral-7B) (d) Micro-level (Vicuna-7B)
Figure 4: Analysis on Graphs with Different $|V|$
@@
Existing graph understanding benchmarks suffer from limited node coverage, unclear task definitions, a lack of structural diversity, and insufficient support for multi-step graph reasoning. To address these issues, we introduce GraphSQA, a comprehensive benchmark designed to evaluate the capabilities of LLMs in understanding graph structures, which encompasses a wide range of task types found in existing benchmarks Wang et al. (2024a); Yuan et al. (2024). GraphSQA includes a broad spectrum of node counts and features diverse graph structures, such as multi-edges and self-loops. It encompasses two types of graph understanding tasks: (1) 5 macro-level graph tasks and (2) 15 micro graph tasks including 7 composite tasks. Further details are provided in Appendix A. In our experiments, we evaluate the average performance of various methods and models on diverse graph tasks with $|V|$ ranging from 15 to 200 nodes, covering 10,400 tasks in total—2,600 macro-level and 7,800 micro-level.
@@
GraphSQA employs three answer types: boolean, numerical, and set. For boolean answers, the score is 1 if correct, otherwise 0. For numerical answers, we use one minus the relative error Wang et al. (2024a). For set answers, the score is the Jaccard similarity Ji et al. (2013) between the answer and the ground truth.
@@
We employed a diverse selection of open-source LLMs, encompassing both long-sequence models, such as Mistral-7B and Llama3-8B-256K (fine-tuned for long sequences), and their counterparts, including Llama3-8B, Qwen2-7B, and Vicuna-7B.
@@ 5.2 Performance
GraphInsight outperforms all other methods across all LLMs for macro-tasks, as shown in Table 4. Specifically, GraphInsight can achieve up to a 4.61$\times$ increase in score compared to AM on Vicuna-7B. Structural methods show the smallest improvement for macro-level tasks. The other two types of methods, i.e., prompting and structural methods, result in only minimal improvements in macro-level tasks. For example, on Llama-3-8B, they can achieve at most a minimal improvement of 5.7%, from 0.4379 to 0.4404. However, GraphInsight can provide a substantial increase of 23.82%, from 0.4379 to 0.5422. Also, as shown in Figures 4(a)-(b), across macro-level tasks with varying $|V|$, GraphInsight consistently outperforms other methods. As $|V|$ increases, the understanding capability tends to decrease, but GraphInsight shows the smallest decline.
@@
(a) Qwen2-7B (b) Vicuna-7B
Figure 5: Analysis on Composite Tasks
(c) Mistral-7B ($\alpha+\beta=15$) (d) Qwen2-7B ($\alpha+\beta=15$)
Figure 6: Head and Tail Strong Region Thresholds $\alpha$ and $\beta$.
(a) Mistral-7B (b) Qwen2-7B
Figure 7: GraphRAG Rate $\gamma\%$.
@@ 5.3 Ablation Study
We conduct the ablation study on graph description organization and GraphRAG to verify the effectiveness of them. As shown in Table 2, graph description organization is the optimization primarily for macro-level tasks as well as for micro-level tasks. For example, on Llama-3-8B-262K, reorganization can achieve an 11.4% increase for macro-level tasks and an 8.9% increase for micro-level tasks. On the other hand, GraphRAG focuses on micro-level tasks and can provide further notably improvements for micro-level tasks on top of the gains achieved by reorganization.
@@ 5.4 Hyperparameter Analysis
We conduct experiments with $\alpha\%+\beta\%=15\%$ (resp. $\frac{\alpha\%}{\beta\%}=\frac{3}{7}$) to assess how variations in $\frac{\alpha\%}{\beta\%}$ (resp. $\alpha\%+\beta\%$) affect LLMs’ graph understanding. As shown in Figure 6, the performance remains stable in both cases, indicating insensitivity to these parameters. We recommend setting $\frac{\alpha\%}{\beta\%}=\frac{3}{7}$ and $\alpha\%+\beta\%=15\%$ for simplicity.
@@
We examine the impact of $\gamma\%$ on the size of the GraphRAG base. As $\gamma\%$ increases, the base retains more critical graph information, improving micro-level understanding. However, as shown in Figure 6, the improvement diminishes when $\gamma\%$ exceeds 80%. We recommend setting $\gamma\%$ to 80%.
@@ 6 Conclusion
In this paper, we introduce GraphInsight, the first framework to address the challenge of graph structure comprehension in LLMs through their inherent “positional bias.” By leveraging strong memory in certain graph sequence regions and compensating for weaker ones, GraphInsight enhances graph understanding. It employs two key techniques: importance-based reorganization and lightweight RAG, optimized for macro- and micro-level tasks. Experiments show that GraphInsight outperforms all baselines across graph sizes. Future work will focus on improving LLMs’ understanding of more complex graph types, including those with labels and semantics.
Limitations
The proposed GraphInsight framework significantly enhances LLMs’ ability to comprehend graph structures, yet it presents several limitations that could inspire future research directions.
First, the quantification of subgraph importance remains a task-dependent challenge, as different graph characteristics and task objectives require distinct importance measures. Future work could explore more task-aware strategies for subgraph importance quantification to improve the framework’s adaptability across various graph-related tasks. Additionally, while the GraphRAG component addresses memory limitations effectively, its optimization in terms of retrieval efficiency and storage overhead remains an open area for improvement. Future research could focus on refining these aspects to enhance retrieval speed and reduce memory consumption. Moreover, the framework’s approach contrasts with other orthogonal LLM-based methods, such as instruction fine-tuning or implicit embedding techniques. Integrating these methods with GraphInsight could complement its capabilities and foster further advancements in graph structure comprehension tasks.
@@ A.1 Introduction to GraphSQA
GraphSQA offers a standardized evaluation suite, enabling consistent and comprehensive assessment of LLMs’ capabilities in understanding and reasoning about graph structures. Unlike existing benchmarks that primarily focus on small graphs containing up to 35 nodes, GraphSQA broadens the scope by encompassing a diverse spectrum of node quantities, ranging from 15 to 300 nodes, and includes various graph structures, such as multi-edges and self-loops. The benchmark is organized into two primary categories of graph understanding tasks: macro-level tasks and micro-level tasks, with the latter further subdivided into 15 tasks, including 8 composite tasks. This comprehensive structure provides a more challenging and extensive evaluation environment, better reflecting the complexities encountered in real-world graph-based applications. A comparative overview of existing benchmarks designed to evaluate the graph understanding capabilities of large language models (LLMs) is presented in Table 3.
@@ A.3 Overview of GraphSQA Categories
  • Triangles Identification: Given the node n, find all triangles (sets of three nodes that are mutually connected) it forms with its neighbors. List the answers in the format of ’[(1, 2, 3), (4, 5, 6), …]’. This task evaluates the model’s ability to identify all triangles that involve a given node and its neighbors. Understanding triangles is fundamental for analyzing local clustering, community structures, and the graph’s overall connectivity.
@@
COT FS BAG BFS DFS SP AL AM
Llama-3-8B-Instruct-262k Overall 3 9 7 4 5 2 5 8 10 1
Macro 3 7 4 8 5 2 6 9 10 1
Micro 3 10 9 2 6 4 5 7 8 1
Meta-Llama-3-8B-Instruct Overall 6 9 5 8 3 2 4 7 10 1
Macro 5 8 6 9 2 3 1 7 10 4
Micro 9 8 5 7 3 2 4 6 10 1
Mistral-7B-Instruct-v0.2 Overall 7 5 9 2 8 6 2 4 10 1
Macro 5 9 2 6 7 4 1 3 10 1
Micro 8 3 9 2 6 5 4 7 10 1
Qwen2-7B-Instruct Overall 5 4 9 7 3 8 6 2 10 1
Macro 7 8 9 6 1 5 3 4 10 1
Micro 3 2 9 7 6 8 4 5 10 1
vicuna-7b-v1.1 Overall 5 5 8 7 4 3 2 9 10 1
Macro 2 3 8 7 5 4 6 9 10 1
Micro 8 6 9 5 4 3 2 7 10 1
@@ Appendix D Notations
Appendix E Experimental Results for different task types
The experimental results for different task types on GraphSQA are illustrated in Figures 8 to 12. These figures analyze macro- and micro-level similarity metrics across various LLM configurations, including Llama-3-8B-Instruct-262k, Meta-Llama-3-8B-Instruct, Mistral-7B-Instruct-v0.2, Qwen2-7B-Instruct, and Vicuna-7b-v1.1.
(a) Macro-level (Llama-3-8B-Instruct-262k)
(b) Micro-level (Llama-3-8B-Instruct-262k)
Figure 8: Macro- and Micro-level Similarity Analysis for Llama-3-8B-Instruct-262k
(a) Macro-level (Meta-Llama-3-8B-Instruct)
(b) Micro-level (Meta-Llama-3-8B-Instruct)
Figure 9: Macro- and Micro-level Similarity Analysis for Meta-Llama-3-8B-Instruct
(a) Macro-level (Mistral-7B-Instruct-v0.2)
(b) Micro-level (Mistral-7B-Instruct-v0.2)
Figure 10: Macro- and Micro-level Similarity Analysis for Mistral-7B-Instruct-v0.2
(a) Macro-level (Qwen2-7B-Instruct)
(b) Micro-level (Qwen2-7B-Instruct)
Figure 11: Macro- and Micro-level Similarity Analysis for Qwen2-7B-Instruct
(a) Macro-level (Vicuna-7b-v1.1)
(b) Micro-level (Vicuna-7b-v1.1)
Figure 12: Macro- and Micro-level Similarity Analysis for Vicuna-7b-v1.1
Appendix F Metric Analysis
@@
Appendix G Performance Analysis of the GraphInsight Framework
@@ Appendix H Discussion
Appendix I Task Templates in GraphSQA